OpenStudy (anonymous):

Lyle purchased a motorcycle for $13,430. It depreciates about 4.1% each year. What is the value of the motorcycle after seven years?

OpenStudy (anonymous):

I started doing this and was on 13,430 +13,430(0.041) = 13,430 +550.03 = 13,980.03

OpenStudy (anonymous):

and continued on

OpenStudy (campbell_st):

use the formula P(1-r)^n \[ A= 13430\times(1-0.041)^7\] you can do the calculation

OpenStudy (anonymous):

i got 12,879.37 but it is not an option

OpenStudy (campbell_st):

I got $10018.60

OpenStudy (anonymous):

how did you get that? I want to learn

OpenStudy (campbell_st):

look at the formula I entered above... I used the compound interest formula... with r as a negative